OpenStudy (anonymous):

can someone explain the integral of 5^x?

OpenStudy (anonymous):

\[\large5^x=e^{\ln(5^x)}=e^{x\ln5}\] \[\text{Let }u=x\ln5\Rightarrow \frac{1}{\ln5}du=dx\]

OpenStudy (anonymous):

Thanks!